OpenStudy (lgbasallote):

\[2(2xy + 4y - 3)dx + (x+2)^2 dy = 0\] how to find the solution to this using linear DE?

OpenStudy (phi):

for an equation in the form \[ \frac{dy}{dx}+P(x)y= Q(x) \] the integrating factor f(x) is \[ f(x)= e^{\int{P(x) dx} }\] and the solution is \[ f(x)y= \int{f(x)Q(x)dx} +c \] It looks like you can put this problem into this form.

OpenStudy (lgbasallote):

how? divide by (x+2)^2 dx?

myininaya (myininaya):

Divide both sides by dx: \[2(2xy+4y-3)+(x+2)^2\frac{dy}{dx}=0\] \[(x+2)^2 \frac{dy}{dx}+4xy+8y-6=0\] Collect y terms together leave y' and y terms on left hand side everything else goes right hand side \[(x+2)^2y'+(4x+8)y=6\] Now to put this equation in the form @phi was talking about yes you need to divide both sides by (x+2)^2

OpenStudy (lgbasallote):

OHHHH!!!! i see it now...thanks so much!!!

OpenStudy (lgbasallote):

so it's linear in x right?

OpenStudy (lgbasallote):

i mean no need for bernoulli's equation yada yada?

myininaya (myininaya):

It is a first order linear differential equation There has actually been a formula derived for these types

OpenStudy (lgbasallote):

really? there is??

myininaya (myininaya):

@phi actually mentions the formula above

myininaya (myininaya):

the part where he says "the solution is ..."

OpenStudy (lgbasallote):

oh that..i thought you had a shortcut or something :p

myininaya (myininaya):

You need to know what P(x) and Q(x) is and find f(x) and just plug in

myininaya (myininaya):

That is the shortcut lol

OpenStudy (lgbasallote):

i thought that was the only way?

myininaya (myininaya):

I can do these long way if I wanted

OpenStudy (lgbasallote):

that means there's a bloodier way o.O

myininaya (myininaya):

ok okay lets talk about the quadratic formula and completing the square what phi gave you was like the "quadratic formula" but we could always "complete the square" instead of using the easy shortcut which is the "quadratic formula"

OpenStudy (lgbasallote):

uhmm could we not :(

myininaya (myininaya):

What?

myininaya (myininaya):

Oh.

OpenStudy (lgbasallote):

just let me believe this is the only way :P lol

myininaya (myininaya):

lol. We don't have to "complete the square" And I hope you know what I mean I'm just trying to give you like analogy

OpenStudy (lgbasallote):

hmm okay..maybe seeing it can help..

myininaya (myininaya):

Ok hey I don't feel like typing it I'm gonna draw it out

myininaya (myininaya):

Is that okay?

OpenStudy (lgbasallote):

sure

myininaya (myininaya):

I mean scan is and junk few moments please

OpenStudy (lgbasallote):

okay :D

OpenStudy (lgbasallote):

why did i uninstall that adobe reader :/ lol

myininaya (myininaya):

Well I think what I just wrote can be found on a lot of different sites anyways

myininaya (myininaya):

Let me see if I can find a good site

OpenStudy (lgbasallote):

okay thanks :DDD

OpenStudy (lgbasallote):

ohh paul's online notes..ive heard rumours bout him

myininaya (myininaya):

So...what i was saying you can go through the steps that actually prove that formula or you can use the formula You know just like if you wanted to use the steps that prove the quadratic formula (completing the square (and there is also my method of course ;) ) ) instead of using the quadratic formula I actually like going the only way on these first order linear differential equations instead of using the formula

myininaya (myininaya):

long* not only sorry

OpenStudy (lgbasallote):

why do you prefer the long way??

myininaya (myininaya):

Because I like the extra work I mean it is like poetry to me when I do it lol I'm weird okay? I prefer the long way on a lot of things like trig substitution and there are other things I just like it Don't know exactly why

OpenStudy (lgbasallote):

lol =))) well i like trig sub too :p i usually prefer step0by-step rather than going by a formula because i dont want memorizing..but this lesson is just something that i should not do step by step lol :))))

myininaya (myininaya):

It is not that bad. lol.

myininaya (myininaya):

Hey one second I think I explained the long way on here before Let me see if I can find it

OpenStudy (lgbasallote):

hmm okay :) i hope it's not that hard...

myininaya (myininaya):

wow I did it the long way here but this was a real confusing one for you at your stage I think i didn't even remember doing this: http://openstudy.com/study#/updates/4e6d116f0b8b4a2b95e1bec4 I think I done an easier one before let me see if I can find it

OpenStudy (lgbasallote):

lol wow you werent kidding @_@

OpenStudy (lgbasallote):

i cant even understand the first step :( lol

myininaya (myininaya):

that might be a little easier because that actually was a first order linear differential equation

OpenStudy (lgbasallote):

i think it resembles bernoulli's equation where you sub v = y^1-n or something like that

OpenStudy (lgbasallote):

i dont think i want this method :( lol

myininaya (myininaya):

Honestly I think I know less math now looking back at this. That is crazy that first link I gave you lol